University Car Wash built a deluxe car wash across the street from campus. The new machines cost $264,000 including installation. The company estimates that the equipment will have a residual value of $25,500. University Car Wash also estimates it will use the machine for six years or about 12,500 total hours. Actual use per year was as follows:

Year Hours Used
1 2,900
2 1,300
3 1,400
4 2,600
5 2,400
6 1,900

I need help preparing the depreciation schedule for six years using the double-depreciation method.

I need help preparing the depreciation schedule for six years using the activity-based method.
Answer #1
1) Depreciation Schedule for six years using the double-depreciation method:
Year Beginning book value Depreciation expense Ending Book Value
a b=a*0.3333 c=a-b
1 $ 2,64,000.00 $ 87,999.99 $ 1,76,000.01
2 $ 1,76,000.01 $ 58,666.66 $ 1,17,333.35
3 $ 1,17,333.35 $ 39,111.11 $     78,222.23
4 $     78,222.23 $ 26,074.08 $     52,148.16
5 $     52,148.16 $ 17,382.72 $     34,765.44
6 $     34,765.44 $ 11,588.48 $     23,176.96
Working:
Straight line depreciation rate = 1/useful life
= 1/6
= 0.16666667
Double declining rate = 2*0.16667
= 0.33333333
2) Year Beginning book value Depreciation expense Ending Book Value
a b c=a-b
1 $ 2,64,000.00 $ 55,332.00 $ 2,08,668.00
2 $ 2,08,668.00 $ 24,804.00 $ 1,83,864.00
3 $ 1,83,864.00 $ 26,712.00 $ 1,57,152.00
4 $ 1,57,152.00 $ 49,608.00 $ 1,07,544.00
5 $ 1,07,544.00 $ 45,792.00 $     61,752.00
6 $     61,752.00 $ 36,252.00 $     25,500.00
Working:
Depreciation rate per hour = (Cost - Salvage Value)/Useful life
= (264000-25500)/12500
= $             19.08
Year Machine usage in hours Depreciation rate per hour Depreciation expense
a b c=a*b
1 2900 $             19.08 $     55,332.00
2 1300 $             19.08 $     24,804.00
3 1400 $             19.08 $     26,712.00
4 2600 $             19.08 $     49,608.00
5 2400 $             19.08 $     45,792.00
6 1900 $             19.08 $     36,252.00